OpenStudy (anonymous):

Ok, so how is cos(2x)=2cos²x-1 ?

OpenStudy (amistre64):

cos(a+b)=cosa cosb - sina sinb if a=b=x cos(x+x)=cos(2x)=cos^2(x) - sin^2(x) can you think of a way to write sin^2 in terms of cos^2?

OpenStudy (anonymous):

well sin²x=1-cos²x right?

OpenStudy (amistre64):

yep, so replace that into the double angle formula and you should be able to simplify it to your original post

OpenStudy (anonymous):

cos(2x)=cos(x+x)=cos(x)cos(x)-sin(x)sin(x)=cos²(x)-sin²(x)=cos²(x)-cos²(x)-1=-1 ??

OpenStudy (amistre64):

c^2-s^2 c^2-(1-c^2) c^2 -1 + c^2 2c^2 -1

OpenStudy (amistre64):

i spose we could have gone with what does -s^2 equal :)

OpenStudy (anonymous):

*sigh*, it's been so long since I needed to do any maths... thanks

OpenStudy (amistre64):

itll be fine, i was outta school for 20+ years before going back for a degree